OpenStudy (anonymous):

if x^2 + px - 5 =0 and x^2 + qx + 5 =0 have a common root the value of (p^2-q^2)=?

OpenStudy (jiteshmeghwal9):

first find the roots of both equations & then the middle term=\[(\alpha+\beta)x\] wht u gt as roots in both of the equation put the values in the expression above

OpenStudy (jiteshmeghwal9):

let the roots of first equation be \[x_1 , y_1\]& the roots of the second equation be\[x_2 , y_2\]write them as follows\[(x_1+y_1)^2-(x_2-y_2)^2\]

OpenStudy (jiteshmeghwal9):

only thinking them as this

OpenStudy (anonymous):

Hey i Think there is some formula if they have CR

OpenStudy (jiteshmeghwal9):

bt dude i m nt sure :/

OpenStudy (jiteshmeghwal9):

@mukushla @experimentX @Callisto m i correct???

OpenStudy (jiteshmeghwal9):

@mathslover m i right????

OpenStudy (anonymous):

@Yahoo! exactly 'one' commen root?

OpenStudy (anonymous):

May Be Two

mathslover (mathslover):

first step must be to find the roots .. of both the eqn ..

mathslover (mathslover):

as mentioned by @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

@mathslover plz first tell me that i m correct or nt ????????????????????

mathslover (mathslover):

@jiteshmeghwal9 please wait u will get your answer by your own soon

OpenStudy (jiteshmeghwal9):

first equations roots\[\alpha={-p + \sqrt{p^2+20}\over2}\]\[\beta={-p-\sqrt{p^2+20}\over2}\]

mathslover (mathslover):

second eqn roots

mathslover (mathslover):

@experimentX what do you think for the next step?

OpenStudy (jiteshmeghwal9):

second equations roots\[\alpha={-q + \sqrt{q^2-20}\over2}\]\[\beta={-q-\sqrt{q^2-20}\over2}\]

OpenStudy (experimentx):

huh??

OpenStudy (jiteshmeghwal9):

\[p={-p+\sqrt{p^2+20}\over2}+{-p-\sqrt{p^2+20}\over2}\]\[q={-q+\sqrt{q^2-20}\over2}-{-q-\sqrt{q^2-20}\over2}\]

OpenStudy (jiteshmeghwal9):

put the values of 'p' in the equation to solve ur question:)

mathslover (mathslover):

@Yahoo! any options given?

OpenStudy (anonymous):

I knw the final answer

mathslover (mathslover):

can u tell it?

OpenStudy (jiteshmeghwal9):

wht is it @Yahoo! ????

OpenStudy (anonymous):

20

OpenStudy (jiteshmeghwal9):

can u plz try mine method & tell me wht u gt?????

mathslover (mathslover):

ok wait i did wrong there. .. posting right one

mathslover (mathslover):

I am getting \(\large{p-q=\sqrt{p^2+20}-\sqrt{q^2-20}}\)

OpenStudy (anonymous):

i want p^2 - q^2?

mathslover (mathslover):

yes wait..

OpenStudy (jiteshmeghwal9):

\[\left( {-p+\sqrt{p^2+20 }\over2}+{-p-\sqrt{p^2+20}\over2} \right)^2-\left( {-q+\sqrt{q^2-20}\over2}+{-q-\sqrt{q^2-20}\over2}= \right)\]=??????????

OpenStudy (jiteshmeghwal9):

anyone can see my reply??????????

OpenStudy (jiteshmeghwal9):

if u solve this u must gt ur answer @Yahoo! :)

OpenStudy (vishweshshrimali5):

Hi everyone well let me tell u all the method to find the common root May be @mukushla will be familiar to it............ We get this expression using method of cross multiplication ......... The method says : Let the common root be a Then put the value of x = a in both eq. and equate them as both of them are roots of the expressions and then solve both eq. using cross multiplication...... This will give u the value of a.....

OpenStudy (experimentx):

i guess 20 ?

OpenStudy (anonymous):

check this... the condition for 2 quadratic equation \(ax^2+bx+c=0\) and \(Ax^2+Bx+C=0\) for have 2 common roots is \(a/A=b/B=c/C\) but for our case this is not true so there is just one root in commen... the condition for 2 quadratic equation \(ax^2+bx+c=0\) and \(Ax^2+Bx+C=0\) for have 1 common roots is \((cA−aC)^2=(bC−cB)(aB−bA)\) it gives \(p^2-q^2=-20\) if im not wrong @Yahoo! try to work it out

OpenStudy (experimentx):